Bibliography

aa

23/03-04 16:20

[andersen1990a] Anders Andersen. Rekesim — simulering av en rekefabrikk. Prosjektrapport, Norsk Institutt for Fiskeri og Havbruksforskning, September 1990.

[andersen1991a] Anders Andersen. En 4-bits adder med ‘fast carry’. Notat D-241, våren 1991, Seksjon for Datafag, Universitetet i Tromsø, February 1991.

[andersen1991b] Anders Andersen. Motorola 68030 ‘pinout’. Notat D-241, våren 1991, Seksjon for Datafag, Universitetet i Tromsø, February 1991.

[andersen1991c] Anders Andersen. Motorola 68030 adresserings modi. Notat D-241, våren 1991, Seksjon for Datafag, Universitetet i Tromsø, March 1991.

[andersen1991d] Anders Andersen. Motorola 680x0 programmering og bruk av debugger/simulator under HP 64000. Notat D-241, våren 1991, Seksjon for Datafag, Universitetet i Tromsø, April 1991.

[andersen1991e] Anders Andersen. En distribuert værapplikasjon for flyplasser. Diplomoppgave, Universitetet i Tromsø, December 1991.

[andersen1992a] Anders Andersen. Bruk av L^{A}TEX til å generere norske dokumenter. Notat, Universitetet i Tromsø, 1992.

[andersen1992b] Anders Andersen. En liten RCS oversikt. Notat, Universitetet i Tromsø, 1992. Kladd.

[andersen1992c] Anders Andersen and Arne Helme. En enkel interpreter i Miranda. Innlvert oppgave d-250, Seksjon for Datafag, Universitetet i Tromsø, May 1992.

[andersen1993a] Anders Andersen. Algebraisk spesifikasjon og korrekt implementasjon av datatypen kø. Notat D-250, våren 1993, Seksjon for Informatikk, Universitetet i Tromsø, April 1993.

[andersen1993b] Anders Andersen. Kø som mutabelt objekt. Notat D-250, våren 1993, Seksjon for Informatikk, Universitetet i Tromsø, April 1993.

[andersen1993c] Anders Andersen. Korrekt implementasjon av stack som mutabelt object. Notat D-250, våren 1993, Seksjon for Informatikk, Universitetet i Tromsø, April 1993.

[andersen1993d] Anders Andersen. TromsTech and the influence from StormCast. Technical report, TromsTech DA, June 1993.

[andersen1993e] Anders Andersen. Using L^{A}TEX at SfI/HDH. Note, Hedmark Distriktshøgskole, November 1993. Local version, Original by Leslie Lamport.

[andersen1994a] Anders Andersen. Hvordan få NetBSD og MS-DOS til å virke i lag på en PC. Notat In226, våren 1994, Hedmark Distriktshøgskole, February 1994.

[andersen1994b] Anders Andersen. Litt om Unix systemprogrammering. Notat In226, våren 1994, Hedmark Distriktshøgskole, March 1994.

[andersen1995a] Anders Andersen. En QoS basert trader for multimediaobjekter i en distribuert omgivelse. Kort beskrivelse av Dr. prosjekt, Universitetet i Tromsø, January 1995.

[andersen1995b] Anders Andersen. Fremdriftsrapport doktorgradstipendiat ELDORADO. Fremdriftsrapport Norges Forskningsråd, Norut IT, July 1995.

[andersen1995c] Anders Andersen. Fremdriftsrapport doktorgradstipendiat informatikk. Fremdriftsrapport Universitetet i Tromsø, Norut IT, September 1995.

[andersen1995d] Anders Andersen. Forvaltning av multimedia tjenestekvalitet. Presentasjon på multimediaseminar høsten 1995, Universitetet i Tromsø, 1995.

[andersen1996a] Anders Andersen. Årsrapport: Doktorstipendiat. Technical report, Norut IT, February 1996.

[andersen1996b] Anders Andersen. Virtuell virkelighet — omdannelse av, forskansing mot eller ny tilgang til virkeligheten. Hjemmeeksamen v301 våren 1996, Universitetet i Tromsø, 1996.

[andersen1996c] Anders Andersen. Application for a visiting research fellowship at distributed multimedia research group in lancaster. Technical report, NORUT IT, June 1996.

[andersen1996d] Anders Andersen. Kommunikasjonsstøtte for multimediaobjekter i en distribuert omgivelse basert på en typemodell. Prosjektsammendrag, NORUT IT, August 1996.

[andersen1996e] Anders Andersen. Forprosjekt om telefoni over Internett. NORUT IT Rapport IT.479/1-96, NORUT IT, December 1996.

[andersen1997a] Anders Andersen. IWQoS’97 og ICODP/ICDP’97. Reiserapport, NORUT Informasjonsteknologi AS, June 1997.

[andersen1997b] Anders Andersen. Doktorgradstipendiat ELDORADO. Fremdriftsrapport, NORUT IT, August 1997.

[andersen1998a] Anders Andersen. A note on reflection in Python 1.5. Distributed Multimedia Research Group Report MPG-98-05, Lancaster University, UK, March 1998.

[andersen1998b] Anders Andersen. Faglig rapport: 101048/420 ELDORADO doktorgradsprogram. Sluttrapport, Norut Informasjonsteknologi AS, March 1998.

[andersen1999a] Anders Andersen. Reflective Middleware and Quality of Service Management. Dr. sci. thesis, Department of Computer Science, University of Tromsø, Tromsø, Norway, February 1999. Draft.

[andersen1999b] Anders Andersen, Gordon S. Blair, Geoff Coulson, and Frank Eliassen. A reflective component-based middleware in Python. Technical report, NORUT IT, 1999. Submitted IPC8.

[andersen1999c] Anders Andersen. The Open-ORB Python prototype API. NORUT IT Report IT302/2-99, NORUT IT, October 1999.

[andersen2000a] Anders Andersen. A reflective component-based middleware in Python. In The Eighth International Python Conference, Arlington, Virginia, USA, January 2000. (short talk).

[andersen2000b] Anders Andersen, Gordon S. Blair, and Frank Eliassen. A reflective componentbased middleware with quality of service management. In PROMS 2000, Protocols for Multimedia Systems, Cracow, Poland, October 2000.

[andersen2000c] Anders Andersen, Gordon S. Blair, and Frank Eliassen. OOPP: A reflective component-based middleware. In NIK 2000, Bodø, Norway, November 2000.

[andersen2001a] Anders Andersen, Gordon S. Blair, Vera Goebel, Randi Karlsen, Tage Stabell-Kulø, and Weihai Yu. Arctic Beans: Configurable and reconfigurable enterprise component architectures. IEEE Distributed Systems Online, 2(7), November 2001. Work in Progress session at Middleware 2001, Heidelberg, Germany.

[andersen2001b] Anders Andersen, Gordon S. Blair, Vera Goebel, Randi Karlsen, Tage Stabell-Kulø, and Weihai Yu. Arctic Beans: Flexible and open enterprise component architectures. In NIK 2001, Tromsø, Norway, November 2001.

[andersen2002a] Anders Andersen. OOPP, A Reflective Middleware Platform including Quality of Service Management. Dr. sci. thesis, Department of Computer Science, University of Tromsø, Tromsø, Norway, February 2002.

[blair1998d] Lynne Blair, Gordon S. Blair, and Anders Andersen. Separating functional behaviour and performance constraints: Aspect-oriented specification. Distributed Multimedia Research Group Report MPG-98-07, Computing Department, Lancaster University, May 1998.

[blair1999a] Gordon S. Blair, Geoff Coulson, Anders Andersen, Lynne Blair, Michael Clarke, Fabio Costa, Hector Duran, Nikos Parlavantzas, and Katia B. Saikoski. A principled approach to supporting adaptation in distributed mobile environments. Technical report, Distributed Multimedia Research Group, Lancaster University, January 1999. (accepte at PDSE’2000, see blair2000a).

[blair1999b] Lynne Blair, Gordon S. Blair, and Anders Andersen. A multi-paradigm specification technique supporting the synthesis of qos management components. Technical report, Distributed Multimedia Research Group, Lancaster University, February 1999.

[blair1999d] Gordon S. Blair, Anders Andersen, Lynne Blair, and Geoff Coulson. The role of reflection in supporting dynamic QoS management functions. In Seventh International Workshop on Quality of Service (IWQoS ’99), number MPG-99-03 in Distributed Multimedia Research Group Report, London, UK, June 1999. IEEE/IFIP, Lancaster University.

[blair1999e] Gordon S. Blair, Anders Andersen, Lynne Blair, Geoff Coulson, and David Sánchez Gancedo. Supporting dynamic QoS management functions in a reflective middleware platform. Technical report, Distributed Multimedia Research Group, Lancaster University, September 1999. (accepted in IEE Proceedings – Software, see blair2000c).

[blair2000a] Gordon S. Blair, Geoff Coulson, Anders Andersen, Lynne Blair, Michael Clarke, Fabio Costa, Hector Duran, Nikos Parlavantzas, and Katia B. Saikoski. A principled approach to supporting adaptation in distributed mobile environments. In 5th International Symposium on Software Engineering for Parallel and Distributed Systems (PDSE-2000), Limerick, Ireland, June 2000.

[blair2000b] Gordon S. Blair, Lynne Blair, Anders Andersen, and Trevor Jones. A formal view of aspects in the development of component-based distributed systems. In SCI’2000 invited session on Generative and Component-Based Software Engineering, Orlando, Florida, USA, July 2000.

[blair2000c] Gordon S. Blair, Anders Andersen, Lynne Blair, Geoff Coulson, and David Sánchez Gancedo. Supporting dynamic QoS management functions in a reflective middleware platform. IEE Proceedings – Software, 147(1):13–21, February 2000.

[blair2001a] Gordon S. Blair, Geoff Coulson, Anders Andersen, Lynne Blair, Michael Clarke, Fabio Costa, Hector Duran-Limon, Tom Fitzpatrick, Lee Johnston, Rui Moreira, Nikos Parlavantzas, and Katia B. Saikoski. The design and implementation of Open ORB 2. IEEE Distributed Systems Online, 2(6), 2001.

[blair2001b] Gordon S. Blair, Lynne Blair, Anders Andersen, and Trevor Jones. Formal support for dynamic QoS management in the development of open component-based distributed systems. IEE Proceedings – Software, 148(3):83–92, June 2001. Special issue on generative and component-based software engineering.

[eliassen1999b] Frank Eliassen, Anders Andersen, Gordon S. Blair, Fabio Costa, Geoff Coulson, Vera Goebel, Øyvind Hanssen, Tom Kristensen, Thomas Plageman, Hans Ole Rafaelsen, Katia B. Saikoski, and Weihai Yu. Next generation middleware: Requirements, architecture, and prototypes. In 7th IEEE Workshop on Future Trends of Distributed Computing Systems (FTDCS ’99), Tunisia, South Africa, December 1999.

[myrvang2001a] Per Harald Myrvang, Tage Stabell-Kulø, and Anders Andersen. Flexible authentication and delegation for distributed components. Technical report, Department of Computer Science, University of Tromsø, 2001. (submitted Middleware 2001).

[papathomas1997b] Michael Papathomas and Anders Andersen. Concurrent object-orieted programming in Python with ATOM. In The Sixth International Python Conference, pages 77–87, San Jose, California, October 1997.